2. Prove the three-set version of the inclusion-exclusion principle: using P(AUB)-P(A) + P(B)
Answer #1

here P(A U B U C) =P((A u B) u C) =P(A u B)+P(C)-P((A u B) n C)

   =P(A)+P(B)-P(A n B)+P(C)-P((A+B - A n B) n C)

=P(A)+P(B)+P(C)-P(A n B)-(P(A n C)+P(B n C)-P(A n B n C))

P(A U B U C)=P(A)+P(B)+P(C)-P(A n B)-P(A n C)-P(B n C)+P(A n B n C)